# Quantifying the Cross-sectoral Intersecting Discrepancies within Multiple Groups Using Latent Class Analysis Towards Fairness

### v3

@@ I Introduction
The Anonymous project [^1] Project name removed to maintain anonymity aims to establish safer online environments for minority ethnic (ME) populations in the UK. Its survey questionnaire covers five key aspects: demography, energy, housing, health, and online services. Notably, the data collected in this context does not directly pertain to resource allocation, making it challenging to explicitly define and detect bias within the data using current methods, despite the presence of discrepancies. These discrepancies may arise from various factors, including culture, user experience, and discrimination, potentially contributing to bias or unfairness. This research is mainly motivated by the Anonymous project, so the datasets we used primarily cover the health, energy, and housing sectors, with our research targeting ME groups.
In our preliminary research (see Appendix A) in England, based on a Anonymous project survey question regarding health and digital services [^2] Question 21: Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?, we observed a notable discrepancy in the Chinese group: 30.16% lacked English proficiency and 26.98% struggled to use the online system, while most other ethnic groups reported less or no concerns. These discrepancies, stemming from cultural differences, user experiences, or discrimination, contribute to inequality and may affect AI fairness. For instance, an AI system might inappropriately assume that only Chinese individuals require English language support, thereby neglecting other ME groups who may also need assistance.
However, investigating multiple and cross-sectoral questions simultaneously is challenging. For instance, the Anonymous project’s data contains similar questions for the energy and housing sectors, and current methods struggle to analyse these sectors jointly. Therefore, we propose an approach to quantify intersecting and cross-sectoral discrepancies for multiple ethnic groups by leveraging latent class analysis (LCA) [17].
LCA is a popular method in social science [18] because it can identify latent groups within a population based on observed characteristics or behaviours. LCA offers a flexible framework for exploring social phenomena and integrating with other analytical techniques. In this research, we use LCA to cluster intersecting and cross-sectoral data, encompassing questions across the health, energy, and housing sectors. This method enables us to derive latent classes and outcomes, with each class describing a distinct cross-sectoral user profile. This approach moves beyond defining user classes solely based on individual questions. More details of our proposed approach are presented in Section II, with experiments and results reported in Section IV.
@@
II Quantifying the Cross-sectoral Intersecting Discrepancies
@@ IV Experiments
IV-A The Anonymous project
Within the Anonymous project444 More dataset details can be found in the Appendix A., a systematic online survey was conducted to explore the experiences of individuals from ME groups in the UK, focusing on digitalised health, energy, and housing aspects. To examine cross-sectoral intersecting discrepancies among the seven ethnic groups (as shown in Table I), we selected three questions from the Anonymous survey data related to these sectors. The answers from ME participants formed the dataset for this experiment. For England and Scotland, we have 594 and 284 samples, respectively. Due to varying sample sizes across ethnicities, we calculated discrepancies separately for each region.
@@ IV-A The Anonymous project
TABLE I: The matrix of discrepancies between 7 ethnic groups for Anonymous project’s England and Scotland data. The AVG denotes the average discrepancy value for one group.
@@ IV-B EVENS
The Centre on the Dynamics of Ethnicity (CoDE), funded by the Economic and Social Research Council (ESRC), conducted “The COVID Race Inequalities Programme”. As part of this project, CoDE carried out the Evidence for Equality National Survey (EVENS) [^5] https://www.evensurvey.co.uk/, which documents the lives of ethnic and religious minorities in Britain during the coronavirus pandemic. The EVENS dataset comprises 14,215 data points and categorizes participants into 18 different ethnic minority groups. To facilitate a comparative analysis with the Anonymous project’s experiment, we focused on the same seven ethnic groups from Anonymous project, resulting in a filtered dataset of 4,348 participants from England and 253 participants from Scotland. We excluded entries with missing values to ensure the robustness of our analysis.
It is important to note that the EVENS dataset uses a different definition for mixed or multiple ethnic groups compared to the Anonymous project. To avoid inconsistencies that could affect the final analysis, we excluded mixed or multiple ethnic groups from the EVENS calculations. Due to the different questions in the EVENS and Anonymous project surveys, we selected three types of cross-sectoral questions from EVENS for analysis: housing, experiences of harassment, and financial situation. These questions were chosen to provide a broad view of participants’ living conditions, social experiences, and economic status during the pandemic.
@@
Seeing the discrepancies in the EVENS England data, shown in the left part of Table II, it is clear that all discrepancy values are relatively small compared to the Anonymous project results. Notably, there are large discrepancies between the Indian and Bangladeshi groups, and between the Indian and Chinese groups. This differs slightly from the conclusions of the Anonymous project’s experiment, likely due to the different types of cross-sectoral questions selected. In the AVG column, the Bangladeshi group has the highest values, indicating they experienced COVID-19 differently. The Census 2021 reported that COVID-19 mortality rates were highest for the Bangladeshi group, for both males and females [32], supporting our findings. The right part of Table II describes the discrepancies for the EVENS Scotland data. There are significant discrepancies between the Caribbean and other ethnic groups, likely related to the unique background of the Caribbean group.
@@ V-B Equal-Size Groups (Quantile-Based)
In Table V, we observed that the discrepancies between Groups 1, 2, 3, 4, and 5 have increased compared to the values presented in Table IV. We believe this is due to all ME LSOAs being concentrated in group 5 under the new grouping method, which explains this observation. In other words, the ME group and non-ME groups have differences in their user profiles. Meanwhile, the largest discrepancy remains between group 1 (predominantly non-ME) and Group 5 (predominantly ME) across the two sampled datasets (70% and 80%) and the original dataset (100%). In Table V, “LR” and “MLP” indicate that the discrepancy calculations were performed on independently sampling datasets. The datasets are used to observe the correlations between discrepancy and AI bias.
@@ VI Conclusion and Limitations
Appendix A The Anonymous project
As part of the Anonymous project, a multilingual online survey, available in 10 languages, was conducted to investigate the experiences of individuals from minority ethnic groups with digitalised housing, health, and energy services. The survey contains a total of 32 questions. The survey data includes 594 responses from England and 284 responses from Scotland. In terms of respondent selection, researcher carefully determined the required number of participants from each ethnic group in England using a proportional allocation method based on their respective population percentages from the 2021 Census (England & Wales). However, due to the unavailability of Scotland’s 2021 census results during the planning phase, the project aimed to limit the number of respondents from each ME group in Scotland to a maximum of 40. The total number of survey respondents was 878. A detailed breakdown of the respondents’ demographic information, including their ethnicities, is provided in Table A.1.
@@ Appendix A The Anonymous project
Anonymous project’s Target Ethnic Group England Scotland Total
@@
We show the distribution of responses from seven ethnic groups in England regarding health and digital services based on Anonymous project’s data. We observed a notable discrepancy in the Chinese group, with 30.16% lacking English proficiency and 26.98% struggling to use the online system, while most other ethnic groups reported no concerns, as shown in Fig. A.1.
Figure A.1: This figure gives an example of the exploratory data analysis (EDA) for the Anonymous project’s England data focuses on Question 21: “Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?” The bars represent the portion of respondents from each ethnic group selecting each option.
Figure A.2: The example of hyperparameter optimisation process to seek the elbow point.
@@
Regarding the selection of the number of latent classes, we used grid search for hyperparameter optimisation with 10-fold cross-validation. The search space for all experiments was adjusted based on empirical observations. In the Anonymous project experiments, we set the range from 2 to 10, while for EVENs and Census experiments, it was set from 2 to 30. One example result of hyperparameter optimisation is shown in Fig. A.2.

### v4

@@ I Introduction
The PRIME project aims to establish safer online environments for minority ethnic (ME) populations in the UK. Its survey questionnaire covers five key aspects: demography, energy, housing, health, and online services. Notably, the data collected in this context does not directly pertain to resource allocation, making it challenging to explicitly define and detect bias within the data using current methods, despite the presence of discrepancies. These discrepancies may arise from various factors, including culture, user experience, and discrimination, potentially contributing to bias or unfairness. This research is mainly motivated by the PRIME project, so the datasets we used primarily cover the health, energy, and housing sectors, with our research targeting ME groups.
In our preliminary research (see Appendix A) in England, based on a PRIME project survey question regarding health and digital services [^1] Question 21: Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?, we observed a notable discrepancy in the Chinese group: 30.16% lacked English proficiency and 26.98% struggled to use the online system, while most other ethnic groups reported less or no concerns. These discrepancies, stemming from cultural differences, user experiences, or discrimination, contribute to inequality and may affect AI fairness. For instance, an AI system might inappropriately assume that only Chinese individuals require English language support, thereby neglecting other ME groups who may also need assistance.
However, investigating multiple and cross-sectoral questions simultaneously is challenging. For instance, the PRIME project’s data contains similar questions for the energy and housing sectors, and current methods struggle to analyse these sectors jointly. Therefore, we propose an approach to quantify intersecting and cross-sectoral discrepancies for multiple ethnic groups by leveraging latent class analysis (LCA) [17].
LCA is a popular method in social science [18] because it can identify latent groups within a population based on observed characteristics or behaviours. LCA offers a flexible framework for exploring social phenomena and integrating with other analytical techniques. In this research, we use LCA to cluster intersecting and cross-sectoral data, encompassing questions across the health, energy, and housing sectors. This method enables us to derive latent classes and outcomes, with each class describing a distinct cross-sectoral user profile. This approach moves beyond defining user classes solely based on individual questions. More details of our proposed approach are presented in Section III, with experiments and results reported in Section IV.
@@
II Related Work
III Quantifying the Cross-sectoral Intersecting Discrepancies
@@ IV Experiments
IV-A The PRIME project
Within the PRIME project333 More dataset details can be found in the Appendix A., a systematic online survey was conducted to explore the experiences of individuals from ME groups in the UK, focusing on digitalised health, energy, and housing aspects. To examine cross-sectoral intersecting discrepancies among the seven ethnic groups (as shown in Table I), we selected three questions from the PRIME survey data related to these sectors. The answers from ME participants formed the dataset for this experiment. For England and Scotland, we have 594 and 284 samples, respectively. Due to varying sample sizes across ethnicities, we calculated discrepancies separately for each region.
@@ IV-A The PRIME project
TABLE I: The matrix of discrepancies between 7 ethnic groups for PRIME project’s England and Scotland data. The AVG denotes the average discrepancy value for one group.
@@ IV-B EVENS
The Centre on the Dynamics of Ethnicity (CoDE), funded by the Economic and Social Research Council (ESRC), conducted “The COVID Race Inequalities Programme”. As part of this project, CoDE carried out the Evidence for Equality National Survey (EVENS) [^4] https://www.evensurvey.co.uk/, which documents the lives of ethnic and religious minorities in Britain during the coronavirus pandemic. The EVENS dataset comprises 14,215 data points and categorizes participants into 18 different ethnic minority groups. To facilitate a comparative analysis with the PRIME project’s experiment, we focused on the same seven ethnic groups from PRIME project, resulting in a filtered dataset of 4,348 participants from England and 253 participants from Scotland. We excluded entries with missing values to ensure the robustness of our analysis.
It is important to note that the EVENS dataset uses a different definition for mixed or multiple ethnic groups compared to the PRIME project. To avoid inconsistencies that could affect the final analysis, we excluded mixed or multiple ethnic groups from the EVENS calculations. Due to the different questions in the EVENS and PRIME project surveys, we selected three types of cross-sectoral questions from EVENS for analysis: housing, experiences of harassment, and financial situation. These questions were chosen to provide a broad view of participants’ living conditions, social experiences, and economic status during the pandemic.
@@
Seeing the discrepancies in the EVENS England data, shown in the left part of Table II, it is clear that all discrepancy values are relatively small compared to the PRIME project results. Notably, there are large discrepancies between the Indian and Bangladeshi groups, and between the Indian and Chinese groups. This differs slightly from the conclusions of the PRIME project’s experiment, likely due to the different types of cross-sectoral questions selected. In the AVG column, the Bangladeshi group has the highest values, indicating they experienced COVID-19 differently. The Census 2021 reported that COVID-19 mortality rates were highest for the Bangladeshi group, for both males and females [32], supporting our findings. The right part of Table II describes the discrepancies for the EVENS Scotland data. There are significant discrepancies between the Caribbean and other ethnic groups, likely related to the unique background of the Caribbean group.
@@ V-B Equal-Size Groups (Quantile-Based)
In Table V, we observed that the discrepancies between Groups 1, 2, 3, 4, and 5 have increased compared to the values presented in Table IV. We believe this is due to all ME LSOAs being concentrated in group 5 under the new grouping method, which explains this observation. In other words, the ME group and non-ME groups have differences in their user profiles. Meanwhile, the largest discrepancy remains between group 1 (predominantly non-ME) and Group 5 (predominantly ME) across the two sampled datasets (90% and 80%) and the original dataset (100%). In Table V, “LR” and “MLP” indicate that the discrepancy calculations were performed on independently sampling datasets. The datasets are used to observe the correlations between discrepancy and AI bias.
@@ VI Conclusion and Limitations
Acknowledgement
This work was supported by URKI’s Engineering and Physical Sciences Research Council (EPSRC) grant numbers EP/W03235X/1, EP/W032333/1, EP/W032341/1, EP/ W032058/1, EP/W032082/1 under the Protecting Minority Ethnic Communities Online (PRIME) project. The authors are also grateful to the PRIME team.
Appendix A The PRIME project
As part of the PRIME project, a multilingual online survey, available in 10 languages, was conducted to investigate the experiences of individuals from minority ethnic groups with digitalised housing, health, and energy services. The survey contains a total of 32 questions. The survey data includes 594 responses from England and 284 responses from Scotland. In terms of respondent selection, researcher carefully determined the required number of participants from each ethnic group in England using a proportional allocation method based on their respective population percentages from the 2021 Census (England & Wales). However, due to the unavailability of Scotland’s 2021 census results during the planning phase, the project aimed to limit the number of respondents from each ME group in Scotland to a maximum of 40. The total number of survey respondents was 878. A detailed breakdown of the respondents’ demographic information, including their ethnicities, is provided in Table A.1.
@@ Appendix A The PRIME project
PRIME project’s Target Ethnic Group England Scotland Total
@@
We show the distribution of responses from seven ethnic groups in England regarding health and digital services based on PRIME project’s data. We observed a notable discrepancy in the Chinese group, with 30.16% lacking English proficiency and 26.98% struggling to use the online system, while most other ethnic groups reported no concerns, as shown in Fig. A.2.
Figure A.1: The example of hyperparameter optimisation process to seek the elbow point. Best viewed when zoomed in.
Figure A.2: This figure gives an example of the exploratory data analysis (EDA) for the PRIME project’s England data focuses on Question 21: “Which of the following concerns do you have about communicating with your general practice (GP) through apps, websites, or other online services?” The bars represent the portion of respondents from each ethnic group selecting each option. Best viewed when zoomed in.
@@
Regarding the selection of the number of latent classes, we used grid search for hyperparameter optimisation with 10-fold cross-validation. The search space for all experiments was adjusted based on empirical observations. In the PRIME project experiments, we set the range from 2 to 10, while for EVENs and Census experiments, it was set from 2 to 30. One example result of hyperparameter optimisation is shown in Fig. A.1.